# Learning Unified Force and Position Control for Legged Loco-Manipulation

### v1

@@ Abstract
Robotic loco-manipulation tasks often involve contact-rich interactions with the environment, requiring the joint modeling of contact force and robot position. However, recent visuomotor policies often focus solely on learning position or force control, overlooking their co-learning. In this work, we propose the first unified policy for legged robots that jointly models force and position control learned without reliance on force sensors. By simulating diverse combinations of position and force commands alongside external disturbance forces, we use reinforcement learning to learn a policy that estimates forces from historical robot states and compensates for them through position and velocity adjustments. This policy enables a wide range of manipulation behaviors under varying force and position inputs, including position tracking, force application, force tracking, and compliant interactions. Furthermore, we demonstrate that the learned policy enhances trajectory-based imitation learning pipelines by incorporating essential contact information through its force estimation module, achieving approximately $\sim$39.5higher success rates across four challenging contact-rich manipulation tasks compared to position-control policies. Extensive experiments on both a quadrupedal manipulator and a humanoid robot validate the versatility and robustness of the proposed policy across diverse scenarios.
Figure 1: We present a unified force-position policy for legged robots that enables diverse loco-manipulation behaviors, including position tracking, force application, and compliant interactions (top). When used for imitation learning data collection, the policy’s learned internal force estimator provides force-aware demonstrations, improving model performance in contact-rich tasks without relying on force sensors (middle). Results on both quadruped and humanoid robots demonstrate the policy’s versatility and robustness (bottom).
@@ 1 Introduction
    We develop a force-aware robot imitation learning data collection pipeline using our learned force estimator, improving position-based imitation learning baselines by $\sim$39.5on three challenging contact-rich manipulation tasks, highlighting our policy’s promise as a general and efficient framework for contact-rich task demonstration curation.
@@ 3.1 A Unified Formulation for Force and Position Control
where ${\bm{F}}_{\text{base}}={\bm{F}}_{\text{base}}^{\text{ext}}+({\bm{F}}_{\text{base}}^{\text{cmd}}-{\bm{F}}_{\text{base}}^{\text{react}})$ is the net force and transform Eq. 2 into:
@@ 3.2 Learning a Unified Force-Position Control Policy
where the input command ${\mathbf{c}}^{\text{cmd}}_{t}=[{\mathbf{v}}_{\text{base}}^{\text{cmd}},{\mathbf{x}}_{\text{ee}}^{\text{cmd}},{\bm{F}}_{\text{ee}}^{\text{cmd}},{\bm{F}}_{\text{base}}^{\text{cmd}}]$ covers the base velocity, end-effector position, end-effector force, and base force commands. For quadrupedal robots, we consider all four command types. For humanoid robots, only the locomotion command ${\mathbf{v}}^{\text{cmd}}_{\text{base}}$ and base force command ${\bm{F}}^{\text{cmd}}_{\text{base}}$ are considered as there is no gripper available on the robot for manipulation tasks. The action ${\mathbf{a}}_{t}$ specifies joint position targets for a PD controller, calculated as ${\mathbf{q}}^{\text{target}}_{t}=\sigma_{a}{\mathbf{a}}_{t}+{\mathbf{q}}^{\text{default}}$, where $\sigma_{a}$ scales the policy output and ${\mathbf{q}}^{\text{default}}$ represents a standard pose.
@@ 4.2 Force-aware Imitation Learning
In Fig. 4 , we compare our method to the baseline in the four real-world tasks. Our approach achieves $\sim$39.5higher success rates than the baseline. In wipe-blackboard, the position-only policy fails to maintain stable contact, often resulting in insufficient wiping or excessive force that risks surface damage. In contrast, our force-aware policy ensures consistent contact pressure, while the low-level policy improves compliance and reduces mechanical stress. For open- and close-cabinet, the primary challenge lies in the push-to-open mechanism’s narrow 3mm stroke, which is difficult to detect using vision alone. Our force estimator accurately senses the required contact force, enabling reliable activation. In open-drawer-occlusion, the baseline policy, relying solely on visual cues, suffers a sharp success rate drop to 0.3 due to unobservable contact. Our method leverages force sensing to detect contact under occlusion, increasing the success rate to 0.76 and underscoring the importance of force estimation in vision-compromised scenarios. We provide all quantitative comparisons and additional details in Appendix D.
@@ Appendix B Hardware Settings and Teleoperation System
As shown in  Fig. 6(b), for manipulation tasks that require only position control, we utilize the MuJoCo AR application, which allows flexible teleoperation of the robot using an iPhone. However, as illustrated in  Fig. 6(a), this approach is no longer suitable for manipulation tasks involving hybrid force and position control. To address this, we developed a dedicated teleoperation system based on the B2 robot’s built-in wireless controller. In this system, two joysticks are used to control the robot’s base movement, while eight buttons below are mapped to control the end-effector’s position and the opening/closing of the gripper. Additionally, by holding the “L1” button, the system switches from issuing position commands to force commands for the end-effector.
@@ C.1 Input Commands and Disturbance Forces
    End-effector force command in cartesian coordinates within the body frame ${\bm{F}}^{\text{cmd}}_{\text{ee}}\in\mathbb{R}^{3}$: $[-60N,60N]$.
@@
[…]
@@
    Base force command within the body frame ${\bm{F}}^{\text{cmd}}_{\text{base}}\in\mathbb{R}^{3}$: $[-60N,60N]$.
@@
    External net force from the environment at the end-effector ${\bm{F}}_{\text{ee}}\in\mathbb{R}^{3}$: $[-60N,60N]$ and at robot base ${\bm{F}}_{\text{base}}\in\mathbb{R}^{3}$: $[-60N,60N]$.
@@ C.2 Reward and Domain Radomization
collision penalty $\mathds{1}_{\text{collision}}$ $-5.0$
joint limit $\mathds{1}_{q>0.8*q^{max}||q<0.8*q^{min}}$ $-10.0$
[…]
action rate $|a_{t-1}-a_{t}|$ $-0.02$
torque limit $\mathds{1}_{\mathbf{\tau}>0.9*|\mathbf{\tau}^{max}|}$ $-0.005$
@@
Friction - [0.3, 2.0]0.32.0[0.3,\,2.0][ 0.3 , 2.0 ]
Body Mass $kg$ [0.0, 15.0]0.015.0[0.0,\,15.0][ 0.0 , 15.0 ]
base com (x,y,z axis) m [0.15, 0.15]0.150.15[-0.15,\,0.15][ - 0.15 , 0.15 ]
Motor Strength % [85, 115]85115[85,\,115][ 85 , 115 ]
Gripper Payload $kg$ [0.0, 0.5]0.00.5[0.0,\,0.5][ 0.0 , 0.5 ]
Push robot $m/s$ [0.0, 0.8]0.00.8[0.0,\,0.8][ 0.0 , 0.8 ], interval = 8s
C.3 Additional Analyses on Policy Learning

### v2

# Learning a Unified Policy for Position and Force Control in Legged Loco-Manipulation

@@ Abstract
Robotic loco-manipulation often involves contact-rich interactions with the environment, requiring the joint modeling of contact force and robot position. However, recent visuomotor policies often focus solely on learning position or force control, overlooking their co-learning. We propose the first unified policy for legged robots that jointly models force and position control learned without relying on force sensors. By simulating diverse combinations of position and force commands alongside external disturbance forces, we use reinforcement learning to learn a policy that estimates forces from historical robot states and compensates for them through position and velocity adjustments. This policy enables a wide range of manipulation behaviors under varying force and position inputs, including position tracking, force application, force tracking, and compliant interactions. Moreover, we demonstrate that the learned policy enhances trajectory-based imitation learning pipelines by incorporating essential contact information through its force estimation module, achieving approximately $\sim$39.5% higher success rates in four challenging contact-rich manipulation tasks over position-control policies. Experiments on both a quadrupedal manipulator and a humanoid robot validate the versatility and robustness of the proposed policy in diverse scenarios.
Figure 1: We present a unified force-position policy for legged robots that enables diverse loco-manipulation behaviors, including position tracking, force application, and compliant interactions (top). When used for imitation learning data collection, the policy’s learned internal force estimator provides force-aware demonstrations, improving model performance in contact-rich tasks without external force sensors (middle). Results on quadruped and humanoid robots demonstrate the policy’s versatility and robustness (bottom).
@@ 1 Introduction
    We develop a force-aware robot imitation learning data collection pipeline using our learned force estimator, improving position-based imitation learning baselines by $\sim$39.5% on three challenging contact-rich manipulation tasks, highlighting our policy’s promise as a general and efficient framework for contact-rich task demonstration curation.
@@ 3.1 A Unified Formulation for Force and Position Control
where we omit the position term because global position of the base is not available. Here, ${\bm{F}}_{\text{base}}={\bm{F}}_{\text{base}}^{\text{ext}}+({\bm{F}}_{\text{base}}^{\text{cmd}}-{\bm{F}}_{\text{base}}^{\text{react}})$ is the net force and transform Eq. 2 into:
@@ 3.2 Learning a Unified Force-Position Control Policy
where the input command ${\mathbf{c}}^{\text{cmd}}_{t}=[{\mathbf{v}}_{\text{base}}^{\text{cmd}},{\mathbf{x}}_{\text{ee}}^{\text{cmd}},{\bm{F}}_{\text{ee}}^{\text{cmd}},{\bm{F}}_{\text{base}}^{\text{cmd}}]$ covers the base velocity, end-effector position, end-effector force, and base force commands. For quadrupedal robots, we consider all four command types. For humanoid robots, only the locomotion command ${\mathbf{v}}^{\text{cmd}}_{\text{base}}$ and base force command ${\bm{F}}^{\text{cmd}}_{\text{base}}$ are considered as there is no gripper available on the robot for manipulation tasks. The output action ${\mathbf{a}}_{t}$ is a residual added to a predefined default pose and ${\mathbf{q}}^{\text{target}}_{t}$ are the joint position targets for the PD controller, calculated as ${\mathbf{q}}^{\text{target}}_{t}=\sigma_{a}{\mathbf{a}}_{t}+{\mathbf{q}}^{\text{default}}$, where $\sigma_{a}$ scales the policy output and ${\mathbf{q}}^{\text{default}}$ represents a standard pose.
@@ 4.2 Force-aware Imitation Learning
In Fig. 4 , we compare our method to the baseline in the four real-world tasks. Our approach achieves $\sim$39.5% higher success rates than the baseline. In wipe-blackboard, the position-only policy fails to maintain stable contact, often resulting in insufficient wiping or excessive force that risks surface damage. In contrast, our force-aware policy ensures consistent contact pressure, while the low-level policy improves compliance and reduces mechanical stress. For open- and close-cabinet, the primary challenge lies in the push-to-open mechanism’s narrow 3mm stroke, which is difficult to detect using vision alone. Our force estimator accurately senses the required contact force, enabling reliable activation. In open-drawer-occlusion, the baseline policy, relying solely on visual cues, suffers a sharp success rate drop to 0.3 due to unobservable contact. Our method leverages force sensing to detect contact under occlusion, increasing the success rate to 0.76 and underscoring the importance of force estimation in vision-compromised scenarios. We provide all quantitative comparisons and additional details in Appendix D.
@@ Appendix A Problem Formulation
Simplified impedance model.
In our formulation, the damping and inertia terms are omitted. This simplification is suitable for relatively static or quasi-static tasks (e.g., wiping or tug-of-war), where motion is slow and normal force ensures contact. In these cases, the static model is sufficient. For dynamic or agile skills, higher control frequencies and explicit velocity/acceleration terms will be necessary, which we plan to explore in future work.
Assumption of rigid contact.
Our formulation ${\bm{F}}=K({\mathbf{x}}-{\mathbf{x}}^{\text{des}})$ naturally handles both rigid and compliant objects, as the end effector position depends only on the net force $\bm{F}$} and the desired goal position ${\mathbf{x}}^{\text{des}}$, regardless of the stiffness of the object. For rigid contacts, small displacements generate the required force, whereas for soft objects, larger deformations occur until the force equilibrium is reached.
@@ Appendix B Hardware Settings and Teleoperation System
As shown in  Fig. 6(b), for manipulation tasks that require only position control, we utilize the MuJoCo AR application, which allows flexible teleoperation of the robot using an iPhone. However, as illustrated in  Fig. 6(a), this approach is no longer suitable for manipulation tasks involving hybrid force and position control. To address this, we developed a dedicated teleoperation system based on the B2 robot’s built-in wireless controller. In this system, two joysticks are used to control the robot’s base movement, while eight buttons below are mapped to control the end-effector’s position and the opening/closing of the gripper. Additionally, by holding the “L1” button, the system switches from issuing position commands to force commands for the end-effector. In practice, data collection with this setup was relatively slow, e.g., approximately 20 seconds per wiping trial and 10 seconds per cabinet trial, mainly due to the limited bandwidth of teleoperation. We believe that exoskeleton-based teleoperation with force feedback will offer a more natural and efficient way to collect demonstrations for forceful manipulation, and we plan to explore this direction in future work.
@@ C.1 Input Commands and Disturbance Forces
    End-effector force command in cartesian coordinates within the body frame ${\bm{F}}^{\text{cmd}}_{\text{ee}}\in\mathbb{R}^{3}$: [60N,60N][-60N,60N].
@@
    Base velocity command ${\mathbf{v}}^{\text{cmd}}_{\text{base}}=(v^{\text{cmd}}_{x},v^{\text{cmd}}_{y},\omega^{\text{cmd}}_{z})$, where $v_{x}\in[-0.8,0.8\,\text{m/s}]$, $v_{y}\in[-0.6,0.6\,\text{m/s}]$, $\omega_{z}\in[-0.8,0.8\,\text{rad/s}]$.
@@
    Base force command within the body frame ${\bm{F}}^{\text{cmd}}_{\text{base}}\in\mathbb{R}^{3}$: [60N,60N][-60N,60N].
@@
    External net force from the environment at the end-effector ${\bm{F}}_{\text{ee}}\in\mathbb{R}^{3}$: [60N,60N][-60N,60N] and at robot base ${\bm{F}}_{\text{base}}\in\mathbb{R}^{3}$: [60N,60N][-60N,60N].
@@ C.2 Reward and Domain Radomization
collision penalty $\mathds{1}_{\text{collision}}$ 5.0-5.0
joint limit $\mathds{1}_{q>0.8*q^{max}||q<0.8*q^{min}}$ 10.0-10.0
torques $|{\mathbf{\tau}}|^{2}$ $5\text{\times}{10}^{-6}$
joint velocities $|\dot{q}|^{2}$ $8\text{\times}{10}^{-4}$
joint acceleration $|\ddot{q}|^{2}$ $2\text{\times}{10}^{-7}$
action rate $|a_{t-1}-a_{t}|$ 0.02-0.02
torque limit $\mathds{1}_{\mathbf{\tau}>0.9*|\mathbf{\tau}^{max}|}$ 0.005-0.005
@@
Friction - [0.3, 2.0][0.3,\,2.0]
Body Mass kgkg [0.0, 15.0][0.0,\,15.0]
base com (x,y,z axis) m [0.15, 0.15][-0.15,\,0.15]
Motor Strength % [85, 115][85,\,115]
Gripper Payload kgkg [0.0, 0.5][0.0,\,0.5]
Push robot $m/s$ [0.0, 0.8][0.0,\,0.8], interval = 8s
C.3 World-Aligned End-Effector Position Estimation
Our estimator predicts not only the external force, but also the end-effector position and base linear velocity. Although forward kinematics provides the end-effector position relative to the arm base, it decouples arm control from base posture. In contrast, we estimate the end-effector position in a world-aligned base frame (with fixed height and orientation relative to the base projection). This representation allows, for example, a downward end-effector command to naturally induce robot base leaning near workspace limits, enabling coordinated whole-body behavior without explicit base control. As the end-effector position in this frame cannot be directly measured, we estimate it instead.
C.4 Additional Analyses on Policy Learning
@@
C.5 Impact of Motor Gains
Compared to pure position-based whole-body control, we use smaller Kp/Kd gains. We observed that lower gains, which allow greater overshoot, improve force estimation. However, excessively low values degrade position tracking accuracy.
C.6 Sim-to-real Gap and Force Estimation Robustness
Our force estimator experiences sim2real gaps due to mismatches in motor dynamics and contact modeling. While domain randomization (e.g., varying Kp/Kd gains) helps reduce these gaps, sim2real transfer remains challenging for RL-based policies. Inspired by how humans rely more on tactile feedback than precise force sensing in contact-rich tasks, our IL policy combines estimated forces with visual input to achieve robust performance without requiring high force accuracy. To further reduce the sim2real gap, we plan to fine-tune the estimator with real-world data and apply system identification methods.